OpenStudy (faiqraees):

Determine the value of m such that the equation x³+mx-1=0 x³-3x+m=0 have a common root @Zarkon @zepdrix

OpenStudy (legomyego180):

Im not the best at these, but is this just asking you to solve for m in terms of x?

jhonyy9 (jhonyy9):

i have a different idea - but i need check it - moment please

OpenStudy (legomyego180):

I've got an answer jhonyy, but im not sure its correct. Im going off the fact that a common zero of two polynomials is a root of their difference.

OpenStudy (mayankdevnani):

^^possible only for quadratic equation provided coeff. of x^2 `1`

OpenStudy (faiqraees):

Okay I came up with a method. But i think I am doing some algebraic mistakes. Start by considering x =a and then make two equations. Now come up with value of a in terms of m i.e. a = m+1/m+3 Next make a different equation by multiplying the first with 3 and second with m. And then come up with value of a^3 in terms of m . Compare the second a^3 with the first a by cubing the first a. Value of m will come

jhonyy9 (jhonyy9):

sorry not is right

jhonyy9 (jhonyy9):

my idea

OpenStudy (faiqraees):

I am pretty sure its correct

jhonyy9 (jhonyy9):

yes this make sens - i agree your idea Faiq

OpenStudy (faiqraees):

x = A [1]-[2] Am-1 +3A -m = 0 A(3+m) = m+1 A = m+1/m+3 3[1] +m[2] 3A³+3Am-3+A³m -3Am+m²=0 A³(3+m)+m²-3=0 A³= (3-m²)/3+m A³= A³ (m+1/m+3)³ = (3-m²)/3+m But this doesnt give me m = -2